OpenStudy (anonymous):

For what value of the variable is the expression undefined? 5/c^2+6c

OpenStudy (anonymous):

\[5 \over c^2+6c\]

OpenStudy (anonymous):

if c^2+6c=0 the expression will be undefined

OpenStudy (accessdenied):

In a rational expression, the 'danger zone' is usually the denominator: \( \large \frac{P(x)}{\color{red}{Q(x)}} \) If \(Q(c) = 0\), then \(c\) would cause the denominator to be 0. Division by 0 is considered undefined (for our purposes). So, that \(c\) value is our interest. \( \large \frac{5}{\color{red}{c^2 + 6c}} \) Similar to the general example; if \(c^2 + 6c = 0\), then \(c\) will cause the denominator to be 0.